OpenStudy (anonymous):

Find an equation for the line with slope m=-7/12 and which goes through the point (36,-5) . Write your answer in the y=mx+b form . Do NOT use decimals as part of your answer...express all parts of your equation in integer or reduced fraction form.

Directrix (directrix):

In the y = mx + b form of a line, m is the slope and b is the y-intercept. m = -7/12 in this problem so we have y = (-7/12) x + b. Now, we have to find b.

Directrix (directrix):

(36,-5) is a point on the line. So, (36,-5) has to "satisfy" or make the equation true. So, -5 = (-7/12) * (36) + b The next step is to sove this equation for b.

Directrix (directrix):

What is -7/12 times 36 ?

OpenStudy (anonymous):

-21?

Directrix (directrix):

Yes. So, -5 = (-7/12) * (36) + b -5 = -21 + b b = 16 -->

Directrix (directrix):

From above, we have y = (-7/12) x + b. b = 16 So, what y = (-7/12) x + ?

OpenStudy (anonymous):

y=-7/12x+16

Directrix (directrix):

That is what I got.

Directrix (directrix):

I might write it as y = (-7/12) x + 16 so that nobody thinks that the x is in a denominator. You write as you do in class.

OpenStudy (anonymous):

thanks(:

Directrix (directrix):

Glad to help.